OpenStudy (anonymous):

I give medals!!! Please help, simple 2 step equations!

OpenStudy (anonymous):

3/5 = 1/5x - 2/5

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

ima look

OpenStudy (anonymous):

if you want to get rid of the fractions, multiply the entire equation by 5

OpenStudy (anonymous):

Yes I know that..

OpenStudy (anonymous):

yay have kellie help he/she is great r u a he or she

OpenStudy (anonymous):

i say girl

OpenStudy (anonymous):

3 = 1x - 2

OpenStudy (anonymous):

3/5 = 1/5x - 2/5 (multiply by 5) 3 = x - 2 (add 2 to each side) 3 + 2 = x 5 = x

OpenStudy (anonymous):

Can you explain it to me though? I don't understand how to do it.

OpenStudy (anonymous):

or... 3/5 = 1/5x - 2/5 (add 2/5 to each side) 3/5 + 2/5 = 1/5x 5/5 = 1/5x (divide by 1/5) (5/5)/(1/5) = x 5/5 * 5/1 = x 25/5 = x 5 = x and yes... I am a girl

OpenStudy (anonymous):

you want me to shoe you the multiplying by 5 part ?

OpenStudy (anonymous):

No, I just need to know how to figure it out, like step by step explaination.

OpenStudy (anonymous):

What do I do first to figure it out?

OpenStudy (anonymous):

are you wanting to deal with the fractions, or get rid of the fractions ??

OpenStudy (anonymous):

Deal with fractions, I don't mind them.

OpenStudy (anonymous):

Becuzz I know how to get rid of them too. So yeah.

OpenStudy (anonymous):

if you look at what I posted above, I did it without fractions and with fractions. Check it out and if you have any questions, just ask

OpenStudy (anonymous):

Ugh, fine. I just need an example: lIKE, STEP 1, DO THIUS, THEN STEP TWO..

OpenStudy (anonymous):

Ugh sorry I hate CAPS, lol

OpenStudy (anonymous):

give me a problem and I will explain step by step, or do you just want me to use the same problem ?

OpenStudy (anonymous):

The same problem is fine.

OpenStudy (anonymous):

3/5 = 1/5x - 2/5 STEP ONE : you want to get the x terms on one side, so we need to add 2/5 3/5 + 2/5 = 1/5x - 2/5 + 2/5 (as you can see, the 2/5 cancel each other out on the right side) STEP TWO : combine like terms 5/5 = 1/5x (to get x by itself we have to divide by 1/5) (5/5)/(1/5) = (1/5x)/(1/5) -- 1/5x divided by 1/5 = 1x so we are left with : (5/5)/(1/5) = x STEP THREE : divide the fractions...when dividing fractions, you " flip " what you are dividing by, then multiply (5/5) / (1/5) = x 5/5 * 5/1 = x 25/5 = x 5 = x does that help ?

OpenStudy (anonymous):

Yes it totally does yay! Okay so I gave a medal and I am a fan. :D

OpenStudy (anonymous):

glad I could help :)

OpenStudy (anonymous):

Thanks, bye.